Find the domain and range of each function. (Enter your answers using interval notation.)
(a)
y = 17/2x + 2


Answer by lwsshak3(11628)

Find the domain and range of each function. (Enter your answers using interval notation.)
(a)
y = 17/2x + 2
Domain: all real numbers except x≠0
(-∞,0) U (0,∞)
..
solve for x in terms of y
17/2x=y-2
lcd:2x
17=2xy-4x=x(2y-4
x=17/(2y-4)
2y-4=0
y≠2
Range:all real numbers except x≠2
(-∞,2) U (2,∞)